ok a few questions?
Have to figure out magnitudes on a ritcher scale for earthquakes with the intensities of the foloowing please help any way you can!
first is 10 7.85. I subnot
second is 10 8.25. I subnot
third is 10 9.6 . I subnot
fourth is 10 6.9 . I subnot?

Answer by stanbon(75887)

Have to figure out magnitudes on a ritcher scale for earthquakes with the intensities of the foloowing please help any way you can!
first is 10 7.85. I subnot
magnitude = 10^7.85 = 70794578.44 Io
------------------------------------------
second is 10 8.25. I subnot
mag = 10^8.25 = 177827941.0 Io
-------------------------------------
third is 10 9.6 . I subnot
mag = 10^9.6 = 3981071706 Io
---------------------------------------
fourth is 10 6.9 Io
mag = 10^6.9 = 7943282.347 Io
===============
Cheers,
Stan H.